A.The equilibrium of income is R17000m.The full employment income is R22000m and the marginal to consume is 0.8.By how much should the investment expenditure change to bring national income to full-employment equilibrium?1.R1 000m 2.1 000m 3.4 000m 4.5 000m
B.which of the following statements is incorrect about the introduction of taxation in the Keynesian model? 1.it increase the value of the multiplier 2.it shifts the aggregate expenditure curve upwards. 3 it increase net exports. 4 it swivels the aggregate expenditure curve downwards. C.C=100+0,80Y what is the savings function in the simple Keynesian model? 1.S=-100 + 0,20Y .2S=100+0,20Y. 3.0,20.4 5
1
Expert's answer
2020-03-11T10:15:34-0400

a) Calculate the value of multiplier

"k=1\/1-MPC" =1/(1-0.8)=1/0.2=5

b) "k=\Delta Y/\Delta I"

ΔY - Change in income

ΔI - change in investment expenses

ΔI=ΔY/k=5000/5=1000

Hence the answer is 1 R 1000 m

b) 1

с)Function Savings - Mirror Image of Consumption Function

Hence the answer is 1